OpenStudy (anonymous):

82 imranmeah91 (or anyone else available) My problem is with the right side of the equation 1/Q dQ=1/435.475-t thank you for continued assistance

OpenStudy (anonymous):

treat the fraction as a constant when you integrate so on the right side you would have constant (fraction) times t and then the integral of -t which is -t^2/2...can you take it from there?

OpenStudy (anonymous):

or is the question 1 / (435.475 - t) ?

OpenStudy (anonymous):

taking the integral of the expression. I cannot see my way clear to get the answer given (320) in the denominator. thanks

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}\ \frac{dQ}{q} \ = \int\limits\limits_{}^{}\ \frac{dt}{435.475-t} \]

OpenStudy (anonymous):

is that accurate so far?